# Continuity of asymptotic entropy on wreath products

### v1

@@ 1. Introduction
This notion was introduced by Avez [Ave72], who proved that if $\mu$ is finitely supported and $h(\mu)=0$, then the random walk $(G,\mu)$ has the Liouville property, meaning that every bounded $\mu$-harmonic function $f:G\to\mathbb{R}$ is constant on the cosets of $\langle\mathrm{supp}(\mu)\rangle$. We recall that a function $f:G\to\mathbb{R}$ is called $\mu$-harmonic if $f(g)=\sum_{h\in G}f(gh)\mu(h)$, for all $g\in G$. The entropy criterion of Derriennic [Der80] and Kaimanovich-Vershik [KV83, Theorem 1.1] states that if $H(\mu)<\infty$, then $h(\mu)>0$ if and only $(G,\mu)$ does not have the Liouville property. Recall that $\mu$ is called non-degenerate if $G=\langle\mathrm{supp}(\mu)\rangle_{+}$ (the semigroup generated by $\mathrm{supp}(\mu)$). If $G$ is a non-amenable group, then any non-degenerate probability measure $\mu$ on $G$ does not have the Liouville property [Aze70, Proposition II.1] [Fur73]. Hence, if in addition $H(\mu)<\infty$, then $h(\mu)>0$. Reciprocally, every amenable group admits a non-degenerate symmetric probability measure $\mu$ with the Liouville property [Ros81, Theorem 1.10] [KV83, Theorem 4.4]. In general, one cannot guarantee in this statement that $H(\mu)<\infty$. Indeed, there are amenable groups on which every non-degenerate probability measure $\mu$ with $H(\mu)<\infty$ does not have the Liouville property [Ers04, Theorem 3.1].
@@
The value $h(\mu)$ determines that, asymptotically, the $n$-th instant of the $\mu$-random walk belongs to a subset of $G$ of size roughly $\exp(h(\mu)n)$; see Remark 2.4 for a precise statement. For random walks with a finite first moment on a finitely generated group, another related quantity is the asymptotic drift $\ell(\mu)\coloneqq\lim_{n\to\infty}\mathbb{E}[|w_{n}|]/n$ of the random walk $\{w_{n}\}_{n\geq 0}$ with respect to a word metric ||$\cdot$|| ⋅ |. In many classes of groups, the ratio $h(\mu)/\ell(\mu)$ coincides with the Hausdorff dimension of the hitting measure of the $\mu$-random walk on a geometric boundary of the group $G$. This is the case for free groups [Kai98, Theorem 2.1.3] and hyperbolic groups [LP08, Theorem 3.1]; see [Tan19, DY23] for further families of groups where this equality holds. We also refer the reader to [Ver00] for a more extensive explanation on the connection between the asymptotic entropy of random walks and other dynamical and algebraic properties of the group.
@@
The question of whether the function $\mu\mapsto h(\mu)$ on a given group $G$ is continuous goes back to [KV83, Section 6, Remark 3]; see also [Ers04, Section 6, Item 5] for an explicit formulation of this problem, and [Ers11a], where the first positive results on continuity of asymptotic entropy are proved. One motivation for this question is that there are multiple situations exhibiting discontinuity in asymptotic entropy within some classes of (amenable) groups. The first examples of such discontinuity were given by Kaimanovich [Kai83, Theorems 3.1 and 4.1], who showed that certain locally finite groups – such as the group of finitely supported permutations of a countable set and the wreath product /2($\mathbb{Z}/2\mathbb{Z}\wr(\bigoplus_{\mathbb{N}}\mathbb{Z}/2\mathbb{Z})$ – admit infinitely supported probability measures with finite entropy, and which do not have the Liouville property. By the entropy criterion, such measures have positive asymptotic entropy, which cannot be approximated by the asymptotic entropy of finitely supported measures. The above examples involve infinitely supported measures, but discontinuity can also occur for probability measures supported on a fixed finite symmetric subset of a group. Indeed, such an example is given in [Ers11a, Lemma 4] on the group $\mathbb{Z}/2\mathbb{Z}\wr D_{\infty}$, where $D_{\infty}$ is the infinite dihedral group.
@@ 1.1. Continuity of entropy on wreath products
Our first main result shows that asymptotic entropy is continuous on wreath products $A\wr B\coloneqq\bigoplus_{B}A\rtimes B$ where the base group $B$ is a hyper-FC-central group with sufficiently fast growth. Recall that a group $G$ is called hyper-FC-central if $G$ has no non-trivial quotient on which any non-trivial element has an infinite conjugacy class. Recall also that a finitely generated group $G$ is said to have at least cubic growth if there is a finite symmetric generating set $S$ of $G$ and a constant $C>0$ such that $|(S\cup\{e_{G}\})^{n}|\geq Cn^{3}$ for all $n\geq 1$.
@@
Let $A$ be a countable group and let $B$ be a countable hyper-FC-central group that contains a finitely generated subgroup of at least cubic growth. Let $\mu$ be a non-degenerate probability measure $\mu$ on $A\wr B$ with $H(\mu)<\infty$. Consider a sequence $\{\mu_{k}\}_{k\geq 1}$ of probability measures on $A\wr B$ with $H(\mu_{k})<\infty$ for all $k\geq 1$, such that $\lim_{k\to\infty}\mu_{k}(g)=\mu(g)$ for every $g\in A\wr B$ and $\lim_{k\to\infty}H(\mu_{k})=H(\mu)$. Then, we have $\lim_{k\to\infty}h(\mu_{k})=h(\mu)$.
@@
In particular, Theorem 1.2 applies when $B=\mathbb{Z}^{d}$, for $d\geq 3$. This is the first known result that establishes the continuity of asymptotic entropy for amenable groups that are not hyper-FC-central, and which works for probability measures with an infinite support. Additionally, it is the first result to address arbitrary sequences of finitely supported probability measures on an amenable group. The only prior result in this direction is [Ers11a, Theorem 5], which shows the continuity of asymptotic entropy for $A\wr\mathbb{Z}$ with $A$ finite, under the assumption that the supports of all probability measures $\mu$, $\mu_{k}$, $k\geq 1$, are contained in a fixed finite subset of $A\wr\mathbb{Z}$. Another feature of Theorem 1.2 is that it holds for all non-degenerate probability measures with finite entropy. For all such measures, the Poisson boundary of $A\wr\mathbb{Z}^{d}$ is non-trivial [Ers04, Theorem 3.1], even in cases where no explicit identification of the Poisson boundary is known. This makes this result the first one to establish the continuity of asymptotic entropy in a setting where the Poisson boundary remains unidentified. Theorem 1.2 is a consequence of a more general criterion for the continuity of asymptotic entropy on wreath products, which is stated in Section 7 (see Theorem 7.1). In Subsection 1.3 we sketch the main ideas behind the proof of Theorem 1.2.
@@
The entropy of a random walk on a wreath product $A\wr B$ at time $n$ is closely related to the number of distinct positions visited by the projection to the base group between times 00 and $n$. This, in turn, is linked to the return probability of the induced random walk on the base group (see Subsection 2.2). To state our next result, we introduce the following notation for the probability that a random walk on a group never returns to the identity element.
@@
This result generalizes [Ers11a, Proposition 2], which does not have hypotheses on the growth of $G$, and supposes that there is a fixed finite subset $F\subseteq G$ such that $\mathrm{supp}(\mu_{k})\subseteq F$ for all $k\geq 1$, and that $\mu$ is symmetric. The proof of Theorem 1.4 is presented in Section 4, and it is based on a uniform version of a comparison lemma due to Coulhon and Saloff–Coste [Cou96, Proposition IV.4] [CSC90, Section II] (see also Proposition 4.3).
@@ 1.2. Connection between the continuity of asymptotic entropy and the weak continuity of harmonic measures
The Poisson boundary $(\partial_{\mu}G,\nu)$ of the $\mu$-random walk on $G$ is a probability space endowed with a measurable $G$-action, such that there is an isomorphism between the space of bounded $\mu$-harmonic functions on $G$ and the space of bounded measurable functions on $(\partial_{\mu}G,\nu)$ [Fur71, Theorem 3.1]. In particular, having the Liouville property is equivalent to the triviality of the Poisson boundary. We refer to Subsection 2.1 for the definition of the Poisson boundary, and to [KV83, Kai00] for a more detailed exposition.
@@
Suppose that there is a Polish space $X$ and probability measures $\nu$, $\{\nu_{k}\}_{k\geq 1}$ on $X$ such that $(X,\nu)$ (resp. $(X,\nu_{k})$) is the Poisson boundary of $(G,\mu)$ (resp. $(G,\mu_{k})$ for each $k\geq 1$). Moreover, suppose that $G$ acts by continuous transformations on $X$.
If the sequence of harmonic measures $\{\nu_{k}\}_{k\geq 1}$ converges weakly to $\nu$, then we have $\lim_{k\to\infty}h(\mu_{k})=h(\mu)$. This holds in particular whenever, in addition to the above assumptions, $X$ is compact and admits a unique $\mu$-stationary probability measure.
@@
    $G$ is a Zariski dense discrete subgroup of an algebraic group and $\mu$, $\{\mu_{k}\}_{k\geq 1}$ have finite entropy and a finite logarithmic moment.
@@
    $G$ is a discrete group of isometries of a $\mathrm{CAT}(0)$ space such that the action of $G$ on $X$ is proper and cocompact, $G$ has a rank one element, and the measures $\mu$, $\{\mu_{k}\}_{k\geq 1}$, have a finite first moment.
@@
    $G$ is a discrete group of isometries of a finite dimensional $\mathrm{CAT}(0)$ cube complex, and the measures $\mu$, $\{\mu_{k}\}_{k\geq 1}$ have finite entropy and a finite logarithmic moment.
@@
We refer to Subsection 8.2 for the definitions of the terms appearing in the statement. The case of acylindrically hyperbolic groups is not new: it has already been proved in [Cho24, Theorem F], and also independently by Joshua Frisch and Anna Erschler in unpublished work. In the particular case of hyperbolic groups, the above was proved in [EK13, Theorem 2] and [GMM18, Theorem 2.9] with additional conditions on the moments of the measures. Our approach is different from the ones used in the previous references. The continuity of asymptotic entropy for the remaining cases of Corollary 1.6 is new, even if one only considers step distributions $\mu$, $\{\mu_{k}\}_{k\geq 1}$ that are supported on a fixed finite generating subset of the group. Some concrete new examples of groups covered by our result are $\mathrm{SL}_{d}(\mathbb{Z})$, $d\geq 3$, and discrete subgroups of isometries of an irreducible Cartan-Hadamard manifold $M$, acting on $M$ properly, cocompactly and with a rank one element. A more extensive and detailed recollection of the classes of groups to which Theorem 1.5 applies is given in Subsection 8.2, where each of the items in Corollary 1.6 is proved.
@@
    In order to recover the value of $\mathcal{P}_{n}^{N}(A\wr B)$ from $w_{n}$, we need to determine both the projection to the base group $B$ as and the lamp configuration every $N$ steps.

### v2

@@ 1. Introduction
This notion was introduced by Avez [AVE72], who proved that if $\mu$ is finitely supported and $h(\mu)=0$, then the random walk $(G,\mu)$ has the Liouville property, meaning that every bounded $\mu$-harmonic function $f:G\to\mathbb{R}$ is constant on the cosets of $\langle\mathrm{supp}(\mu)\rangle$. We recall that a function $f:G\to\mathbb{R}$ is called $\mu$-harmonic if $f(g)=\sum_{h\in G}f(gh)\mu(h)$, for all $g\in G$. The entropy criterion of Derriennic [DER80] and Kaimanovich-Vershik [KV83, Theorem 1.1] states that if $H(\mu)<\infty$, then $h(\mu)>0$ if and only if $(G,\mu)$ does not have the Liouville property. Recall that $\mu$ is called non-degenerate if $G=\langle\mathrm{supp}(\mu)\rangle_{+}$ (the semigroup generated by $\mathrm{supp}(\mu)$). If $G$ is a non-amenable group, then any non-degenerate probability measure $\mu$ on $G$ does not have the Liouville property [AZE70, Proposition II.1] [FUR73]. Hence, if in addition $H(\mu)<\infty$, then $h(\mu)>0$. Reciprocally, every amenable group admits a non-degenerate symmetric probability measure $\mu$ with the Liouville property [ROS81, Theorem 1.10] [KV83, Theorem 4.4]. In general, one cannot guarantee in this statement that $H(\mu)<\infty$. Indeed, there are amenable groups on which every non-degenerate probability measure $\mu$ with $H(\mu)<\infty$ does not have the Liouville property [ERS04, Theorem 3.1].
@@
The value $h(\mu)$ determines that, asymptotically, the $n$-th instant of the $\mu$-random walk belongs to a subset of $G$ of size roughly $\exp(h(\mu)n)$; see Remark 2.4 for a precise statement. For random walks with a finite first moment on a finitely generated group, another related quantity is the asymptotic drift $\ell(\mu)\coloneqq\lim_{n\to\infty}\mathbb{E}[|w_{n}|]/n$ of the random walk $\{w_{n}\}_{n\geq 0}$ with respect to a word metric $|\cdot|$. In many classes of groups, the ratio $h(\mu)/\ell(\mu)$ coincides with the Hausdorff dimension of the hitting measure of the $\mu$-random walk on a geometric boundary of the group $G$. This is the case for free groups [KAI98, Theorem 2.1.3] and hyperbolic groups [LE 08, Theorem 3.1]; see [TAN19, DY23] for further families of groups where this equality holds. We also refer the reader to [VER00] for a more extensive explanation on the connection between the asymptotic entropy of random walks and other dynamical and algebraic properties of the group.
@@
The question of whether the function $\mu\mapsto h(\mu)$ on a given group $G$ is continuous was first explicitly asked in [ERS04, Section 6, Item 5], and later the first positive results on continuity of asymptotic entropy were given in [ERS11a]. One motivation for this problem is that there are multiple situations exhibiting discontinuity in asymptotic entropy within some classes of (amenable) groups. The first examples of such discontinuity were given by Kaimanovich [KAI83, Theorems 3.1 and 4.1], who showed that certain locally finite groups – such as the group of finitely supported permutations of a countable set and the wreath product $\mathbb{Z}/2\mathbb{Z}\wr(\bigoplus_{\mathbb{N}}\mathbb{Z}/2\mathbb{Z})$ – admit infinitely supported probability measures with finite entropy, and which do not have the Liouville property; see also [KV83, Section 6, Remark 3]. By the entropy criterion, such measures have positive asymptotic entropy, which cannot be approximated by the asymptotic entropy of finitely supported measures. The above examples involve infinitely supported measures, but discontinuity can also occur for probability measures supported on a fixed finite symmetric subset of a group. Indeed, such an example is given in [ERS11a, Lemma 4] on the group $\mathbb{Z}/2\mathbb{Z}\wr D_{\infty}$, where $D_{\infty}$ is the infinite dihedral group. It is claimed by Vershik in the paragraph following Lemma 2 in [VER00, page 682] that the estimates from [VER99] can be used to show that asymptotic entropy is continuous in the set of non-degenerate symmetric probability measures supported on a fixed finite generating set of a group. Note that Erschler’s example does not contradict Vershik’s claim, since it consists of a sequence of non-symmetric probability measures that converge to a degenerate probability measure. However, we have not been able to reconstruct Vershik’s proof, and we believe that the author might have meant to say “upper-semicontinuous” instead of “continuous”, which is enough to prove [VER00, Lemma 2]. We remark that the results we prove in the current paper guarantee the continuity of asymptotic entropy for possibly non-symmetric and infinitely supported probability measures.
In the literature there are several results on regularity properties of asymptotic entropy that go beyond continuity, such as Lipschitz continuity, differentiability and analyticity. Families of groups studied so far include free groups [LED12], hyperbolic groups [LED13, GOU17, GL13] and acylindrically hyperbolic groups [MS20, Theorem 13.2], for probability measures with suitable moment conditions and that in particular include measures supported on a fixed finite symmetric generating subset of the group. In the current paper we focus exclusively on continuity results, and the study of further regularity properties of asymptotic entropy on wreath products will be developed elsewhere.
@@ 1.1. Continuity of entropy on wreath products
Our first main result shows that asymptotic entropy is continuous on wreath products $A\wr B\coloneqq\bigoplus_{B}A\rtimes B$ where the base group $B$ is a hyper-FC-central group with sufficiently fast growth. Recall that a group $G$ is called hyper-FC-central if $G$ has no non-trivial quotient on which any non-trivial element has an infinite conjugacy class. Recall also that a group $G$ has at least cubic growth if there is a finite symmetric subset $S\subseteq G$ such that $\liminf_{n\to\infty}|S^{n}|/n^{3}>0$. Any finitely generated subgroup of a countable group that does not have at least cubic growth is has a finite index subgroup that is either trivial, $\mathbb{Z}$ or $\mathbb{Z}^{2}$.
@@
Let $A$ be a countable group and let $B$ be a countable hyper-FC-central group of at least cubic growth. Let $\mu$ be a non-degenerate probability measure $\mu$ on $A\wr B$ with $H(\mu)<\infty$. Consider a sequence $\{\mu_{k}\}_{k\geq 1}$ of probability measures on $A\wr B$ with $H(\mu_{k})<\infty$ for all $k\geq 1$, such that $\lim_{k\to\infty}\mu_{k}(g)=\mu(g)$ for every $g\in A\wr B$ and $\lim_{k\to\infty}H(\mu_{k})=H(\mu)$. Then, we have $\lim_{k\to\infty}h(\mu_{k})=h(\mu)$.
For a finitely generated group, being hyper-FC-central is equivalent to having a nilpotent subgroup of finite index [DM56, Theorem 2] [MCL56, Theorem 2] (see also [FF18] for a self-contained proof of this result). Therefore, hyper-FC-central groups of at least cubic growth contain a copy of either $\mathbb{Z}^{3}$ or the Heisenberg group $H_{3}(\mathbb{Z})$. Theorem 1.2 applies in particular when $B=\mathbb{Z}^{d}$, for $d\geq 3$. This is the first known result that establishes the continuity of asymptotic entropy for amenable groups that are not hyper-FC-central, and which works for probability measures with an infinite support. Additionally, it is the first result to address arbitrary sequences of finitely supported probability measures on an amenable group. The only prior result in this direction is [ERS11a, Theorem 5], which shows the continuity of asymptotic entropy for $A\wr\mathbb{Z}$ with $A$ finite, under the assumption that the supports of all probability measures $\mu$, $\mu_{k}$, $k\geq 1$, are contained in a fixed finite subset of $A\wr\mathbb{Z}$. Note that Erschler’s example of discontinuity consider wreath products over $D_{\infty}$, which has linear growth. This shows that the hypothesis of at least cubic growth of $B$ cannot be removed. The hypotheses of the convergence of the Shannon entropies of the measures is also natural, since for infinitely supported probability measures it is possible to have the pointwise convergence $\lim_{k\to\infty}\mu_{k}(g)=\mu(g)$ for each $g\in G$ and still not have the convergence of $\{H(\mu_{k})\}_{k\geq 1}$ to $H(\mu)$. Such examples would also provide probability measures for which asymptotic entropy is not continuous. Note that if one further supposes that all the measures $\{\mu_{k}\}_{k\geq 1}$ are supported on a fixed finite subset of $G$, then the convergence of Shannon entropies is implied by the pointwise convergence of the probability measures. This is the setting considered by Erschler in [ERS11a], called “strong convergence” there. The non-degeneracy assumption on the probability measure $\mu$ in Theorem 1.2 can be replaced with the condition that the semigroup generated by $\mathrm{supp}(\mu)$ is a subgroup of $A\wr B$ whose projection to the base group $B$ has at least cubic growth. We stated Theorem 1.2 with the non-degeneracy assumption for clarity in the exposition of the result.
@@
A feature of Theorem 1.2 is that it holds for all non-degenerate probability measures with finite entropy. For all such measures on $A\wr\mathbb{Z}^{d}$, the Poisson boundary is non-trivial [ERS04, Theorem 3.1]. This makes Theorem 1.2 the first one to establish the continuity of asymptotic entropy in a setting where the Poisson boundary remains unidentified. Theorem 1.2 is a consequence of a more general criterion for the continuity of asymptotic entropy on wreath products, which is stated in Section 7 (see Theorem 7.1). In Subsection 1.3 we sketch the main ideas behind the proof of Theorem 1.2.
@@
The entropy of a random walk on a wreath product $A\wr B$ at time $n$ is closely related to the number of distinct positions visited by the projection to the base group $B$ between times 0 and $n$. This, in turn, is linked to the return probability of the induced random walk on the base group (see Subsection 2.2). To state our next result, we introduce the following notation for the probability that a random walk on a group never returns to the identity element.
@@
This result generalizes [ERS11a, Proposition 2], which does not have hypotheses on the growth of the subgroup generated by $\mathrm{supp}(\mu)$, but supposes that there is a fixed finite subset $F\subseteq G$ such that $\mathrm{supp}(\mu_{k})\subseteq F$ for all $k\geq 1$, and that $\mu$ is symmetric. The proof of Theorem 1.4 is presented in Section 4, and it is based on a uniform version of a comparison lemma due to Coulhon and Saloff–Coste [COU96, Proposition IV.4] [CS90, Section II] (see also Proposition 4.3).
@@ 1.2. Connection between the continuity of asymptotic entropy and the weak continuity of harmonic measures
The Poisson boundary $(\partial_{\mu}G,\nu)$ of the $\mu$-random walk on $G$ is a probability space endowed with a measurable $G$-action, such that there is a $G$-equivariant isomorphism between the space of bounded $\mu$-harmonic functions on $G$ and the space of bounded measurable functions on $(\partial_{\mu}G,\nu)$ [FUR71, Theorem 3.1]. In particular, having the Liouville property is equivalent to the triviality of the Poisson boundary. We refer to Subsection 2.1 for the definition of the Poisson boundary, and to [KV83, KAI00] for a more detailed exposition.
@@
Suppose that there is a Polish space $X$ on which $G$ acts by continuous transformations, and Borel probability measures $\nu$, $\{\nu_{k}\}_{k\geq 1}$ on $X$ such that $(X,\nu)$ (resp. $(X,\nu_{k})$) is the Poisson boundary of $(G,\mu)$ (resp. $(G,\mu_{k})$ for each $k\geq 1$).
If the sequence $\{\nu_{k}\}_{k\geq 1}$ converges weakly to $\nu$, then we have $\lim_{k\to\infty}h(\mu_{k})=h(\mu)$. This holds in particular whenever, in addition to the above assumptions, $X$ is compact and admits a unique $\mu$-stationary probability measure.
@@
    $G$ is a Zariski dense discrete subgroup of $\mathrm{SL}_{d}(\mathbb{R})$ and $\mu$, $\{\mu_{k}\}_{k\geq 1}$ have finite entropy and a finite logarithmic moment.
@@
    $G$ is a subgroup of isometries of a proper $\mathrm{CAT}(0)$ space $X$ such that the action of $G$ on $X$ is proper and cocompact, $G$ has a rank one element, and the measures $\mu$, $\{\mu_{k}\}_{k\geq 1}$, have a finite first moment.
@@
    $G$ is a discrete subgroup of isometries of a finite dimensional $\mathrm{CAT}(0)$ cube complex $X$ such that the action of $G$ on $X$ is proper, and the measures $\mu$, $\{\mu_{k}\}_{k\geq 1}$ have finite entropy and a finite logarithmic moment.
@@
We refer to Subsection 8.2 for the definitions of the terms appearing in the statement. The case of acylindrically hyperbolic groups is not new: it has already been proved in [CHO24, Theorem F], and also independently by Anna Erschler and Joshua Frisch in unpublished work. In the particular case of hyperbolic groups, the above was proved in [EK13, Theorem 2] and [GMM18, Theorem 2.9] with additional conditions on the moments of the measures. Our approach is different from the ones used in the previous references. The continuity of asymptotic entropy for the remaining cases of Corollary 1.6 is new, even if one only considers step distributions $\mu$, $\{\mu_{k}\}_{k\geq 1}$ that are supported on a fixed finite generating subset of the group. Some concrete new examples of groups covered by our result are $\mathrm{SL}_{d}(\mathbb{Z})$, $d\geq 3$, and discrete subgroups of isometries of an irreducible Cartan-Hadamard manifold $M$, acting on $M$ properly, cocompactly and with a rank one element. A more extensive and detailed recollection of the classes of groups to which Theorem 1.5 applies is given in Subsection 8.2, where each of the items in Corollary 1.6 is proved. In particular, we show that a special case of Theorem 1.2, under additional assumptions on the measures $\mu$ and $\mu_{k}$, $k\geq 1$, can be deduced from Theorem 1.5. However, this argument does not recover the result in full generality; see Corollary 8.13 and the paragraph following it.
@@
    In order to recover the value of $\mathcal{P}_{n}^{N}(A\wr B)$ from $w_{n}$, we need to determine both the projection to the base group $B$ and the lamp configuration every $N$ steps.